Evaluate the limit as x turns to 2+

lim {sin[In (x-1)] / √(x-2)}


1
Expert's answer
2021-10-25T16:44:56-0400

"\lim\limits_{x\rightarrow2^+} \frac{sin(ln(x-1))}{\sqrt{x-2}}\\
=\lim\limits_{x\rightarrow2^+} \frac{sin(ln(x-1))}{(ln(x-1))} \times \frac{(ln(x-1))}{\sqrt{x-2}}\\
=\lim\limits_{x\rightarrow2^+} \frac{sin(ln(x-1))}{(ln(x-1))} \times \lim\limits_{x\rightarrow2^+}\frac{(ln(x-1))}{\sqrt{x-2}}\\
=(1)\times \lim\limits_{x\rightarrow2^+}\frac{(ln(x-1))}{\sqrt{x-2}}\\"

Now, applying L'Hopital's rule, we get:

"\lim\limits_{x\rightarrow2^+} \frac{\frac{1}{x-1}}{\frac{1}{2\sqrt{x-2}}}\\
=\lim\limits_{x\rightarrow2^+} \frac{2\sqrt{x-2}}{x-1}\\
=\frac{2(\sqrt{2-2})}{2-1}\\
=0"